Hankins Corporation has 6.9 million shares of common stock outstanding, 250,000 shares of 4.2 percent preferred stock outstanding, par value of $100; and 135,000 bonds with a semiannual coupon rate of 5.7 percent outstanding, par value $1,000 each. The common stock currently sells for $67 per share and has a beta of 1.05, the preferred stock has a par value of $100 and currently sells for $89 per share, and the bonds have 18 years to maturity and sell for 107 percent of par. The market risk premium is 7.7 percent, T-bills are yielding 2.8 percent, and the company’s tax rate is 24 percent.

a.

What is the firm’s market value capital structure? (Do not round intermediate calculations and round your answers to 4 decimal places, e.g., .1616.)

b. If the company is evaluating a new investment project that has the same risk as the firm’s typical project, what rate should the firm use to discount the project’s cash flows? (Do not round intermediate calculations enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

a. Debt

b. Preferred stock

Equity

Discount rate%
Answer #1
Instrument Number Market Value / unit Market Value Capital structure
N P MV = P x N MV / total MV
Debt       135,000                            1,070    144,450,000 0.2297
Preferred stock       250,000                                  89      22,250,000 0.0354
Equity    6,900,000                                  67    462,300,000 0.7350
Total    629,000,000

Part (a)

The last column in the table above contains your answer.

Debt: 0.2297

Preferred stock: 0.0354

Equity: 0.7350

Part (b)

Pre tax cost of debt = YTM = 2 x semi annual YTM = 2 x RATE (Nper, PMT, PV, FV) = RATE (2 x 18, 5.7% / 2 x 1000, -1070, 1000) = 5.10%

Tax rate, T = 24%

Post tax cost of debt = YTM x (1 - T) = 5.10% x (1 - 24%) = 3.88%

Cost of preferred stock = Div / Price = 4.2% x 100 / 89 = 4.72%

Cost of common stock = risk free rate + beta x market risk premium = 2.8% + 1.05 x 7.7% = 10.89%

Discount rate = WACC = 9.06% as calculated below

Instrument Capital structure Post tax cost Component cost
A B A x B
Debt 0.2297 3.88% 0.89%
Preferred stock 0.0354 4.72% 0.17%
Equity 0.7350 10.89% 8.00%
Total 9.06%